OpenStudy (anonymous):

The perimeter of a rectangle with base 10 in. and height 13 in. is _____. Be sure to include units? is it 130?

hartnn (hartnn):

the perimeter is 2(base+height) so 2(10+13) = 23*2 = 46 in. ok ?

OpenStudy (anonymous):

oh ok thanks